# Belief Scene Graphs: Expanding Partial Scenes with Objects through Computation of Expectation

### v1

@@ III BELIEF SCENE GRAPHS
This section details the design process of the novel etablished CECI model, as well as the construction of the dataset for the training process. In addition, it presents the methodology for the generation of a Belief Scene Graph.
@@ III-A CECI Model Architecture Design and Optimization
Depending on the diversity and scale of the environments, the number of nodes $\mathcal{V}$ may be substantial. In order to tackle this, the set of nodes $\mathcal{V}$ is masked to include only building and room nodes, which represent our set of nodes of interest $\mathcal{N}$. This is possible since the object nodes are already encompassed as node attributes in the set $\mathcal{N}$ (i.e. $\textbf{E}^{\prime}_{\mathcal{N}}(C)\times O^{\prime}_{\mathcal{N}}$). By doing so, we can optimize the learning process and the overall performance of the network. Under this considerations we propose a novel GCN architecture, namely CECI model. The propose architecture consist of 5 GCN convolutional layers [16], followed by batch normalization [17], ReLU and dropout. The input to the network is a given graph $\mathcal{G}^{\prime}$ with a set of histograms $\textbf{E}_{\mathcal{N}}^{\prime}(C)\times O^{\prime}_{\mathcal{N}}$ as node attribute and the output is the predicted set of histograms $\textbf{E}_{\mathcal{N}}^{\prime\prime}(C)\times O^{\prime\prime}_{\mathcal{N}}$, both with size $Q\times L$, where $Q\in\mathbb{Z}^{+}$ is the number of building and room nodes and $L\in\mathbb{Z}^{+}$ is the number of labeled object classes. The overall network architecture can be visualize in Fig. 1.
@@ III-B Dataset Generation for CECI Model Training
Due to the novelty of the concept, currently there are not available datasets for the training of the CECI model. In order to generate the data needed for training, we used the Habitat-Matterport 3D Research Dataset (HM3D) [18], currently the world’s largest dataset of real-life residential, commercial, and civic spaces. First, we generated a custom-made mapping involving the grouping and filtering of the 1659 semantic categories present on the dataset, into a subset of 45 hand-picked object class labels selected based on their frequency on the scenes and their relevance to a robotic task. In addition, we have chosen to delimit the original set of room labels to a single general label (i.e. ”room”) [19] to allow the proper generalization of the framework to unstructured indoor environments. This is an important consideration since the method is meant to work in partially known environments, where the semantic labels of the rooms are not always reliable. Then, for each of the dataset’s semantically annotated 3D spaces, we constructed a 3D scene graph that includes 3 layers (i.e. building, rooms and objects) and the nodes consist of one attribute (i.e. the class label), while the edges represent descendant relationships (i.e. the child node is physically contained on the parent node).
For each ground truth graph $\mathcal{G}$, we generated in addition a series of input graphs $\mathcal{G}^{\prime}$ by deleting object nodes at random, until the number of deleted nodes is equal to 50% of the original number of nodes. The training data generation process was designed to resemble, as close as possible, the information that a typical autonomous robotic system would acquire when deployed in a real-life mission and presented with different perception challenges (e.g. occluded objects, out-of-view objects, not visited areas, etc). Then the information of the remaining object nodes is encoded as a room node attribute, by encompassing the object nodes into a histogram of the 45 labeled object classes $C$ over the respective room node. Where the histogram represents the probability distribution $\textbf{E}_{\mathcal{R}}^{\prime}(C)\times O^{\prime}_{\mathcal{R}}$ over the room node $\mathcal{R}\in\mathcal{N}$, for the defined set of objects $C$, where $O^{\prime}_{\mathcal{R}}\in\mathbb{Z}^{+}$ is the number of object nodes connected to the room. Likewise, the ground truth histogram $\textbf{E}_{\mathcal{R}}(C)\times O_{\mathcal{R}}$ is estimated in the same way. In addition, the histograms $\textbf{E}_{\mathcal{H}}(C)\times O_{\mathcal{H}}$ and $\textbf{E}_{\mathcal{H}}^{\prime}(C)\times O_{\mathcal{H}}^{\prime}$ for the building node $\mathcal{H}\in\mathcal{N}$ of every pair $\mathcal{G}$ and $\mathcal{G}^{\prime}$ is also computed. The resulting dataset consisted of 4,525 graph pairs with a total of 582,000 instances of the 45 labeled object classes divided across 65,225 rooms.
@@ III-C Incorporating Expectation on Belief Scene Graphs
The Belief Scene Graph is defined as a tuple of nodes and edges $\mathcal{G}^{\prime\prime}=(\{\mathcal{B,\mathcal{V}}\},\mathcal{E})$, where $\mathcal{E}$ is the set of directional edges, $\mathcal{V}$ is the set of nodes presented in the original 3D scene graph $\mathcal{G}^{\prime}$, with the set of predicted histograms $\textbf{E}_{\mathcal{N}}^{\prime\prime}(C)\times O_{\mathcal{N}}^{\prime\prime}$, as node attribute for the building and room nodes, and $\mathcal{B}$ is the set of blind nodes. In a broad sense, a Belief Scene Graph could be seen as an expansion of a 3D scene graph, through the incorporation of the concept of expectation. In this work, we represent expectation in two main ways, the first one is in the shape of the expected probabilistic distribution over a room node of the possible object nodes connected to it, the second one, is using a new kind of nodes called Blind Nodes, which represent the expected objects to be found.
These nodes are generated in different ways, for example, we can complete the graph using the difference between the histograms of the input 3D scene graph $\textbf{E}_{\mathcal{R}}^{\prime}(C)\times O^{\prime}_{\mathcal{R}}$ and the predicted one $\textbf{E}_{\mathcal{R}}^{\prime\prime}(C)\times O^{\prime\prime}_{\mathcal{R}}$ by adding connections from the room nodes to Blind Nodes with the respective classes until the number of nodes and the class distribution of the room nodes matches the predicted one. On the other hand, we can also add only the Blind Nodes of the labeled classes with the higher expectation corresponding to objects that were not originally present in the input 3D scene graph, but based on the current information of the environment are expected or believed to be there. The layer and information of these expectations depends on the specific task at hand, for example, in an exploration scenario, we might want to know the probabilities for certain zones to be found on unexplored paths, while in a pick-and-drop scenario, we may rather estimate the size and number of objects present in a set of pick-up locations.
@@ IV EXPERIMENTAL EVALUATION
In this section, we first provide the training parameters of the CECI model, followed by the validation using Wasserstein distance [20], energy distance [21], Mean Error (ME) and Frobenius norm [22], in addition, a qualitative analysis is also presented. Secondly, we present the simulation results for single-object and multi-object search. Finally, we report the results for prediction of unseen-objects in a real-world indoor environment with a Boston Dynamics Spot Legged Robot.
@@ IV-B Validation Metrics
During the validation, three main metrics were taken into consideration, the first one was the statistical distance between the probability distributions of the predicted histogram and the ground truth. For this we computed the Wasserstein distance (i.e. the earth mover’s distance) [20] and the energy distance [21] for the pair of probabilistic distributions $\textbf{E}_{\mathcal{N}}^{\prime\prime}(C)$ for the prediction and $\textbf{E}_{\mathcal{N}}(C)$ for the ground truth. Table I shows the mean, variance, skewness and kurtosis for each of the computed distances. Overall, it can be observed that the predicted distribution is fairy similar to the ground truth since it presents a low mean in both metrics in addition to a low variance.
@@
Finally, we estimated the Frobenius norm [22] of the difference between the correlation matrices of the predicted and the ground truth data, the computed value 6.95 indicates a fair degree of similarity and showcase the performance of the CECI model for learning the underlying correlations between the different object class labels present in the generated detest. Fig. 3 presents both correlation matrices, where the main differences can be observed on the objects that are room-agnostic (can be found in multiple rooms, e.g. curtain and bottle). This behaviour is consistent with Fig. 2.
@@ IV-C Qualitative Analysis
The comparison between the input, predicted and ground truth histogram in Fig. 3(a) helps to illustrate the overall performance of the proposed model, in which the bigger differences are present at the object class labels with the higher frequency on the scene (e.g. clothes and floor mat), as well as in the ones with frequency equal to 0 (e.g. bookshelf and wardrobe). More importantly, the capability of the model to predict objects that were not present in the input graph can be observed, for instance toilet and desk. This capability in particular is crucial, and showcases the merit of the CECI method to compute expectation based on the correlation information of the graph taxonomy, which is one of the main contributions of this work.
The computed histogram can then be used to complete the 3D scene graph with Blind Nodes as previously mentioned in III-C, as well as to extract the probability distribution $\textbf{E}_{\mathcal{N}}^{\prime\prime}(C)$ of the respective building or room node. Fig. 3(b) shows the distributions for the histograms presented in Fig. 3(a), which helps to illustrate the results presented in Table I. This is a key property to tackle the problem of object search efficiently, while different nodes may contain an instance of the artifact, the distribution of the room allow us to prioritize the ones which are more likely to contain it, in order to optimize the search base on the most likely node candidate. In other words, the one with the highest expectation of containing the artifact.
@@ IV-D Belief Scene Graphs for Object Search Task
For the specifics of the experimental evaluation presented in this article, the application scenario will be centered around an object search task [24, 25, 26] in simulation. The objective is to find an artifact or artifacts (i.e. an specific object class) in a partially known environment. This task was selected based on the high level of complexity for a variety of real-life robotic missions, like rescue operations, active source seeking or multi-agent task planning.
Object Search Definition: Given a partial representation of the environment, represented by a given graph $\mathcal{G}^{\prime}=(\mathcal{V},\mathcal{M})$, where $\mathcal{M}$ is the set of directional edges $e_{ij}(v_{i},v_{j}):(\mathcal{V}\times\mathcal{V})\mapsto\mathcal{M}$ representing the cost of traversing from the node $v_{i}$ to the node $v_{j}$. In the task of object search, we need to find the optimal policy $\mathcal{P}^{*}$, which comprises of the set of edges to be traversed to reach the node $\mathcal{S}\in\mathcal{V}$ that contains the artifact $\alpha\notin\mathcal{V}$. For this task the robotic system considers a graph $\mathcal{G}^{\prime}$ with traversal cost as edge attribute as follows:
  • Building layer. The evaluation scenario only considers single building environments.
@@
For comparison purposes, two algorithms were implemented, one using a 3D scene graph and other using the proposed Belief Scene Graph. More specifically, Dijkstra’s graph search approach [27] has been selected for the 3D Scene Graph, where the planner optimizes the traversal of the environment by prioritizing the shortest path to an unexplored room until the artifact is found. Regarding the Belief Scene Graph, the graph search planner optimizes the traversal of the environment by prioritizing the room with the highest expectation of containing the artifact as $max(\textbf{E}_{\mathcal{R}}^{\prime\prime}(\alpha))\mapsto e_{ij}$.
Using the test split of the dataset, both algorithms were evaluated in a series of object search scenarios. The simulation runs were designed such that the position of the robot in the environment is randomized in every run, under the condition that it can not start in a room where the artifact is present. In addition, $e$ was randomly assigned for each simulated environment of the test split. The artifact is also switched to a different class label every certain number of runs. As such, the objective is for the robotic system to find an artifact or artifacts in an partially known environment with the lowest traversal cost. Once a room/area has been explored, the graph $\mathcal{G}^{\prime}$ will be updated with the rest of the objects present in that room. The task is finished once an instance of the artifact has been found for the first case scenario or once all the instances of the artifact are found for the second case scenario.
The simulation results showed that the total traversal cost computed over multiple randomized runs for the single-artifact case and the multi-artifact case is on average 33.8% and 45.5% lower, respectively, when using $\mathcal{G}^{\prime\prime}$ compared to $\mathcal{G}^{\prime}$. Demonstrating that the use of BSG based planners yield better result overall for the task of object search compared to the classical 3DSG method.
@@ V CONCLUSIONS
A novel concept of Belief Scene Graphs was introduced in this work, as a utility driven extension of the classical 3D scene graphs. Reasoning about presence/absence of new objects through expectations represents a crucial incremental step in the way robotics systems understand their environment, enabling close-to-humans task planing and task optimization. The proposed CECI model (histograms learnt through GCN training) enabled the translation of an intuitive representation of beliefs into a computational method. The output histograms of the CECI model were then used to generate a Belief Scene Graph by completing the 3D scene graph with Blind Nodes and to extract the probabilistic distributions of the expected objects. The overall framework was experimentally validated for the task of object search. In addition we present a field experiment in a real-world indoor environment showcasing the ability of the method to level with human common sense. Future work will focus on the implementation of the framework on a broad range of robotic applications like exploration and task allocation for multi-agent scenarios.

### v2

@@ III BELIEF SCENE GRAPHS
This section details the design process of the novel established CECI model, as well as the construction of the dataset for the training process. In addition, it presents the methodology for the generation of a Belief Scene Graph.
@@ III-A CECI Model Architecture Design and Optimization
Depending on the diversity and scale of the environments, the number of nodes $\mathcal{V}$ may be substantial. In order to tackle this, the set of nodes $\mathcal{V}$ is masked to include only building and room nodes, which represent our set of nodes of interest $\mathcal{N}\subset\mathcal{V}$. This is possible since the object nodes are already encompassed as node attributes in the set $\mathcal{N}$ (i.e. $E^{\prime}_{\mathcal{N}}(C)\times O^{\prime}_{\mathcal{N}}$). By doing so, we can optimize the learning process and the overall performance of the network. Under these considerations, we propose a novel GCN architecture, namely CECI model. The proposed architecture consists of 5 GCN convolutional layers [16], followed by batch normalization [17], ReLU and dropout. The input to the network is a given graph $\mathcal{G}^{\prime}$ with a set of histograms $E_{\mathcal{N}}^{\prime}(C)\times O^{\prime}_{\mathcal{N}}$ as node attribute and the output is the predicted set of histograms $E_{\mathcal{N}}^{\prime\prime}(C)\times O^{\prime\prime}_{\mathcal{N}}$, both with size $q\times n$, where $q\in\mathbb{Z}^{+}$ is the number of building and room nodes and $n\in\mathbb{Z}^{+}$ is the number of labeled object classes. The overall network architecture can be visualized in Fig. 1.
@@ III-B Dataset Generation for CECI Model Training
Due to the novelty of the concept, currently there are no available datasets for the training of the CECI model. In order to generate the data needed for training, we used the Habitat-Matterport 3D Research Dataset (HM3D) [18], currently the world’s largest dataset of real-life residential, commercial, and civic spaces. First, we generated a custom-made mapping involving the grouping and filtering of the 1659 semantic categories present on the dataset, into a subset of 45 hand-picked object class labels selected based on their frequency on the scenes and their relevance to a robotic task. In addition, we have chosen to delimit the original set of room labels to a single general label (i.e. ”room”) [19, 13, 14] to allow the proper generalization of the framework to unstructured indoor environments. This is an important consideration since the method is meant to work in partially known environments, where the semantic labels of the rooms are not always reliable. Then, for each of the dataset’s semantically annotated 3D spaces, we constructed a 3D scene graph that includes 3 layers (i.e. building, rooms and objects) and the nodes consist of one attribute (i.e. the class label), while the edges represent descendant relationships (i.e. the child node is physically contained on the parent node).
For each ground truth graph $\mathcal{G}$, we generated in addition a series of input graphs $\mathcal{G}^{\prime}$ by deleting object nodes at random, until the number of deleted nodes is equal to 50% of the original number of nodes. The training data generation process was designed to resemble, as closely as possible, the information that a typical autonomous robotic system would acquire when deployed in a real-life mission and presented with different perception challenges (e.g. occluded objects, out-of-view objects, not visited areas, etc). Then the information of the remaining object nodes is encoded as a room node attribute, by encompassing the object nodes into a histogram of the 45 labeled object classes $C$ over the respective room node. Where the histogram represents the probability distribution $E_{\rho}^{\prime}(C)\times O^{\prime}_{\rho}$ over the room node $\rho\in\mathcal{N}$, for the defined set of objects $C$, where $O^{\prime}_{\rho}\in\mathbb{Z}^{+}$ is the number of object nodes connected to the room. Likewise, the ground truth histogram $E_{\rho}(C)\times O_{\rho}$ is estimated in the same way. In addition, the histograms $E_{\eta}(C)\times O_{\eta}$ and $E_{\eta}^{\prime}(C)\times O_{\eta}^{\prime}$ for the building node $\eta\in\mathcal{N}$ of every pair $\mathcal{G}$ and $\mathcal{G}^{\prime}$ is also computed.
@@ III-C Incorporating Expectation on Belief Scene Graphs
The Belief Scene Graph is defined as a tuple of nodes and edges $\mathcal{G}^{\prime\prime}=(\{\mathcal{B,\mathcal{V}}\},\mathcal{E})$, where $\mathcal{E}$ is the set of directional edges, $\mathcal{V}$ is the set of nodes presented in the original 3D scene graph $\mathcal{G}^{\prime}$, with the set of predicted histograms $E_{\mathcal{N}}^{\prime\prime}(C)\times O_{\mathcal{N}}^{\prime\prime}$, as node attribute for the building and room nodes, and $\mathcal{B}$ is the set of blind nodes. In a broad sense, a Belief Scene Graph could be seen as an expansion of a 3D scene graph, through the incorporation of the concept of expectation. In this work, we represent expectation in two main ways, the first one is in the shape of the expected probabilistic distribution over a room node of the possible object nodes connected to it, and the second one is using a new kind of nodes called blind nodes, which represent the expected objects to be found.
These nodes are generated in different ways, for example, we can complete the graph using the difference between the histograms of the input 3D scene graph $E_{\rho}^{\prime}(C)\times O^{\prime}_{\rho}$ and the predicted one $E_{\rho}^{\prime\prime}(C)\times O^{\prime\prime}_{\rho}$ by adding connections from the room nodes to blind nodes with the respective classes until the number of nodes and the class distribution of the room nodes matches the predicted one. On the other hand, we can also add only the blind nodes of the labeled classes with the higher expectation corresponding to objects that were not originally present in the input 3D scene graph, but based on the current information of the environment are expected or believed to be there. The layer and information of these expectations depend on the specific task at hand, for example, in an exploration scenario, we might want to know the probabilities for certain zones to be found on unexplored paths, while in a pick-and-drop scenario, we may rather estimate the size and number of objects present in a set of pick-up locations.
@@ IV EXPERIMENTAL EVALUATION
In this section, we first provide the training parameters of the CECI model, followed by the validation using Wasserstein distance [20], energy distance [21], Mean Error (ME) and Frobenius norm [22], in addition, a qualitative analysis is also presented. Secondly, we present the simulation results for single-object and multi-object search. Finally, we report the results for the prediction of unseen-objects in a real-world indoor environment with a Boston Dynamics Spot Legged Robot.
@@ IV-B Validation Metrics
During the validation, three main metrics were taken into consideration, the first one was the statistical distance between the probability distributions of the predicted histogram and the ground truth. For this we computed the Wasserstein distance (i.e. the earth mover’s distance) [20] and the energy distance [21] for the pair of probabilistic distributions $E_{\mathcal{N}}^{\prime\prime}(C)$ for the prediction and $E_{\mathcal{N}}(C)$ for the ground truth. Table I shows the mean, variance, skewness and kurtosis for each of the computed distances. Overall, it can be observed that the predicted distribution is fairly similar to the ground truth since it presents a low mean in both metrics in addition to a low variance.
@@
Finally, we estimated the Frobenius norm [22] of the difference between the correlation matrices of the predicted and the ground truth data. The computed value 6.95 indicates a fair degree of similarity and showcases the performance of the CECI model for learning the underlying correlations between the different object class labels present in the generated data set. Fig. 3 presents both correlation matrices, where the main differences can be observed on the objects that are room-agnostic (i.e. can be found in multiple rooms, e.g. curtain and bottle). This behavior is consistent with Fig. 2.
@@ IV-C Qualitative Analysis
The comparison between the input, predicted and ground truth histogram in Fig. 4(a) helps to illustrate the overall performance of the proposed model, in which the bigger differences are present at the object class labels with the higher frequency on the scene (e.g. clothes and floor mat), as well as in the ones with a frequency equal to 0 (e.g. bookshelf and wardrobe). More importantly, the capability of the model to predict objects that were not present in the input graph can be observed, for instance toilet and desk. This capability in particular is crucial and showcases the merit of the CECI method to compute expectation based on the correlation information of the graph taxonomy, which is one of the main contributions of this work.
The computed histogram can then be used to complete the 3D scene graph with blind nodes as previously mentioned in III-C, as well as to extract the probability distribution $E_{\mathcal{N}}^{\prime\prime}(C)$ of the respective building or room node. Fig. 4(b) shows the distributions for the histograms presented in Fig. 4(a), which helps to illustrate the results presented in Table I. This is a key property to tackle the problem of object search efficiently, while different nodes may contain an instance of the artifact, the distribution of the room allows us to prioritize the ones that are more likely to contain it, in order to optimize the search base on the most likely node candidate. In other words, the one with the highest expectation of containing the artifact.
@@ IV-D Belief Scene Graphs for Object Search Task
For the specifics of the experimental evaluation presented in this article, the application scenario will be centered around an object search task [24, 25, 26] in simulation. The objective is to find an artifact or artifacts (i.e. a specific object class) in a partially known environment. This task was selected based on the high level of complexity for a variety of real-life robotic missions, like rescue operations, active source seeking or multi-agent task planning.
Object Search Definition: Given a partial representation of the environment, represented by a given graph $\mathcal{G}^{\prime}=(\mathcal{V},\mathcal{E})$, where $\mathcal{E}$ is the set of directional edges $e_{ij}(v_{i},v_{j}):(\mathcal{V}\times\mathcal{V})\mapsto\mathcal{M}$ representing the cost of traversing from the node $v_{i}$ to the node $v_{j}$. In the task of object search, we need to find the optimal policy $\mathcal{P}^{*}$, which comprises of the set of edges to be traversed to reach the node $\mathcal{S}\in\mathcal{V}$ that contains the artifact $\alpha\notin\mathcal{V}$. For this task, the robotic system considers a graph $\mathcal{G}^{\prime}$ with traversal cost as edge attribute as follows:
  • Building layer. The evaluation scenario only considers single-building environments.
@@
For comparison purposes, two algorithms were implemented, one using a 3D scene graph and the other using the proposed Belief Scene Graph. More specifically, Dijkstra’s graph search approach [27] has been selected for the 3D Scene Graph, where the planner optimizes the traversal of the environment by prioritizing the shortest path to an unexplored room until the artifact is found. Regarding the Belief Scene Graph, the graph search planner optimizes the traversal of the environment by prioritizing the room with the highest expectation of containing the artifact as $max(E_{\rho}^{\prime\prime}(\alpha))\mapsto e_{ij}$.
Using the test split of the dataset, both algorithms were evaluated in a series of object search scenarios. The simulation runs were designed such that the position of the robot in the environment is randomized in every run, under the condition that it can not start in a room where the artifact is present. In addition, $e$ was randomly assigned for each simulated environment of the test split. The artifact is also switched to a different class label every certain number of runs. As such, the objective is for the robotic system to find an artifact or artifacts in a partially known environment with the lowest traversal cost. Once a room/area has been explored, the graph $\mathcal{G}^{\prime}$ will be updated with the rest of the objects present in that room. The task is finished once an instance of the artifact has been found for the first case scenario or once all the instances of the artifact are found for the second case scenario.
The simulation results showed that the total traversal cost computed over multiple randomized runs for the single-artifact case and the multi-artifact case is on average 33.8% and 45.5% lower, respectively, when using $\mathcal{G}^{\prime\prime}$ compared to $\mathcal{G}^{\prime}$. Demonstrating that the use of BSG-based planners yields better results overall for the task of object search compared to the classical 3DSG method.
@@ IV-E Field Test Results for CECI
Additionally, to the simulation comparisons of the object search task presented in Section IV-D, we tested our approach on the Boston Dynamics Spot legged robot in a real civic indoor environment with the goal to contrast Belief Scene Graph expectation with human common sense for unseen-objects. The environment consists of a small waiting room (i.e. Room A), a shared kitchen with a dining area (i.e. Room B), and a meeting room (i.e. Room C). The objective was for the robot to traverse and generate a 3D scene graph with partial information about the objects present in the scene, this graph was then enhanced using the proposed CECI model to draw expectations about other possible objects in the environment that were not part of the ones seen during the experiment.
@@ V CONCLUSIONS
A novel concept of Belief Scene Graphs was introduced in this work, as a utility-driven extension of the classical 3D scene graphs. Reasoning about the presence/absence of new objects through expectations represents a crucial incremental step in the way robotics systems understand their environment, enabling close-to-human task planning and task optimization. The proposed CECI model (histograms learnt through GCN training) enabled the translation of an intuitive representation of beliefs into a computational method. The output histograms of the CECI model were then used to generate a Belief Scene Graph by completing the 3D scene graph with blind nodes and to extract the probabilistic distributions of the expected objects. The overall framework was experimentally validated for the task of object search. In addition, we present a field experiment in a real-world indoor environment showcasing the ability of the method to level with human common sense. Future work will focus on the implementation of the framework on a broad range of robotic applications like exploration and task allocation for multi-agent scenarios.